# Statistical learnability of smooth boundaries via pairwise binary classification with deep ReLU networks

### v1

@@ A.1 Proof of Proposition 3
Note that $\eta(x,x^{\prime})=0$ otherwise. Thus, the distribution satisfies the conditions (A2) and (A4). Therefore, the distribution of $(X,X^{\prime},Y)$, denoted by $P$, belongs to $\mathcal{P}_{\alpha}$. By the definition, we have $[\mathscr{S}]=\mathscr{S}_{P}$. ∎
@@ A.7.2 Step 2
Let $\alpha>0$, and let $\mathcal{C}^{\alpha}([-2^{-1},2^{-1}]^{d_{0}-1})$ be the unit ball of Hölder space defined on the domain $[-2^{-1},2^{-1}]^{d_{0}-1}$. Let $\mathcal{K}\subset[-2^{-1},2^{-1}]^{d_{0}}$ be any set such that there are function $h\in\mathcal{C}^{\alpha}([-2^{-1},2^{-1}]^{d_{0}-1})$ and index $i\in\{1,\cdots,d_{0}\}$ satisfying
@@
Let $0<\varepsilon<2^{-1}$. Let $P\in\mathcal{P}_{\alpha}$, and let $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$ be any sequence in $\mathscr{S}_{P}$. Denote by $g_{i}^{*}=\mathds{1}_{\mathcal{K}_{i}}$, $i=1,\cdots,d_{1}$. Then, there are ReLU networks $g_{\bm{W}^{*},\bm{b}^{*}}\in\mathcal{F}_{L^{*},J^{*},S^{*},M^{*},\bm{d}^{*}}^{\textup{NN}}(\mathcal{X})$ with constant $c\geq 0$ and parameters $L^{*}\lesssim\log_{2}\varepsilon^{-1}+1$, $J^{*}\lesssim\varepsilon^{-c}\vee 1$, $S^{*}\lesssim\varepsilon^{-(d_{0}-1)/\alpha}(1\vee\log_{2}\varepsilon^{-1})$, $M^{*}\lesssim|\log(4\varepsilon)|\vee 1$, and $\bm{d}^{*}=(d_{0},d_{\textup{NN},1}^{*},\cdots,d_{\textup{NN},L^{*}-1}^{*},d_{1})$ such that for every $i=1,\cdots,d_{1}$,
@@ A.7.4 Step 4
To apply the results in (Park, , 2009; Kim et al., , 2021), we need to evaluate the covering numbers of the class of classifiers. Given a compact subset $\mathcal{A}$ in a measure space equipped with the Lebesgue measure, let $\mathscr{N}(\delta,\mathcal{H},\|\cdot\|_{L^{\infty}(\mathcal{A})})$ denote the covering number of a class $\mathcal{H}$ of real-valued functions with respect to the pseudo-distance $\|\cdot\|_{L^{\infty}(\mathcal{A})}$ (see e.g., (Steinwart and Christmann, , 2008, Definition 6.19) for covering numbers). Besides, given a map $h:\Delta^{d}\to\mathcal{A}$ for a set $\mathcal{A}$, denote by $h\circ\mathcal{F}=\{h\circ f\;|\;f\in\mathcal{F}\}$. Then, the following evaluation holds.
@@
$\log\mathscr{N}(\delta,\psi\circ\rho\circ\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime}),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
@@
$\mathscr{N}(\delta,\psi\circ\rho\circ\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime}),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
$\leq\mathscr{N}(2^{-1}D_{\Delta^{d}}^{2}\delta,\rho\circ\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime}),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})}).$ (50)
@@
$\mathscr{N}(2^{-1}D_{\Delta^{d}}^{2}\delta,\rho\circ\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime}),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
@@
$\mathscr{N}(\delta,\psi\circ\rho\circ\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime}),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
@@
$\log\mathscr{N}(\delta,\psi\circ\rho\circ\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime}),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
@@
Note that the subclass $\mathcal{F}^{*}=\mathcal{F}_{L^{*},J^{*},S^{*},M^{*},\bm{d}^{*}}^{\Delta^{d}\textup{-NN}}$ is considered in this proof. Denote by $\mathcal{F}_{\textup{local}}^{*}=\mathscr{F}_{\beta,\beta^{-1}\varepsilon_{n},P}(\mathcal{F}^{*})$. Then, the set $\psi\circ\rho\circ\mathcal{F}_{\textup{local}}^{*}$ is $1$-uniformly bounded. Also, it is well known that hinge loss is $1$-Lipschitz continuous in the first argument (see, e.g., (Steinwart and Christmann, , 2008, Example 2.27)). Thus, we can set $c_{0}=1$ and $M_{n}=1$ hereafter.
@@
$\log\mathscr{B}(s,\psi\circ\rho\circ\mathcal{F}_{\textup{local}}^{*},\|\cdot\|_{L^{2}(\mathcal{X}^{2})})\leq\log\mathscr{N}(\frac{s}{2},\psi\circ\rho\circ\mathcal{F}_{\textup{local}}^{*},\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})}).$ (63)
@@
$\log\mathscr{N}(2^{-1}s,\psi\circ\rho\circ\mathcal{F}_{\textup{local}}^{*},\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$

### v2

@@ A.1 Proof of Proposition 3
Note that $\eta(x,x^{\prime})=0$ otherwise. Thus, the distribution satisfies the conditions (A2) and (A4). Therefore, the distribution of $(X,X^{\prime},Y)$, denoted by $P$, belongs to $\mathcal{P}_{\alpha}$. By the definition, we have $\mathscr{S}\sim\mathscr{S}_{P}$. ∎
@@ A.7.2 Step 2
Let $\alpha>0$, $R\geq 0$, and let $\mathcal{C}^{\alpha}([-2^{-1},2^{-1}]^{d_{0}-1})$ be the ball of Hölder space defined on the domain $[-2^{-1},2^{-1}]^{d_{0}-1}$ with radius $R$. Let $\mathcal{K}\subset[-2^{-1},2^{-1}]^{d_{0}}$ be any set such that there are function $h\in\mathcal{C}^{\alpha}([-2^{-1},2^{-1}]^{d_{0}-1})$ and index $i\in\{1,\cdots,d_{0}\}$ satisfying
@@
Let $0<\varepsilon<2^{-1}$. Let $P\in\mathcal{P}_{\alpha}$, and let $\mathscr{S}_{P}=\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$. Denote by $g_{i}^{*}=\mathds{1}_{\mathcal{K}_{i}}$, $i=1,\cdots,d_{1}$. Then, there are ReLU networks $g_{\bm{W}^{*},\bm{b}^{*}}\in\mathcal{F}_{L^{*},J^{*},S^{*},M^{*},\bm{d}^{*}}^{\textup{NN}}(\mathcal{X})$ with constant $c\geq 0$ and parameters $L^{*}\lesssim\log_{2}\varepsilon^{-1}+1$, $J^{*}\lesssim\varepsilon^{-c}\vee 1$, $S^{*}\lesssim\varepsilon^{-(d_{0}-1)/\alpha}(1\vee\log_{2}\varepsilon^{-1})$, $M^{*}\lesssim|\log(4\varepsilon)|\vee 1$, and $\bm{d}^{*}=(d_{0},d_{\textup{NN},1}^{*},\cdots,d_{\textup{NN},L^{*}-1}^{*},d_{1})$ such that for every $i=1,\cdots,d_{1}$,
@@ A.7.4 Step 4
To apply the results in (Park, , 2009; Kim et al., , 2021), we need to evaluate the covering numbers of the class of classifiers. Given a compact subset $\mathcal{A}$ in a measure space equipped with the Lebesgue measure, let $\mathscr{N}(\delta,\mathcal{H},\|\cdot\|_{L^{\infty}(\mathcal{A})})$ denote the covering number of a class $\mathcal{H}$ of real-valued functions with respect to the pseudo-distance $\|\cdot\|_{L^{\infty}(\mathcal{A})}$ (see e.g., (Steinwart and Christmann, , 2008, Definition 6.19) for covering numbers). Also, for any $\mathcal{F}\subset\mathcal{F}_{0}$, denote by $\rho(\mathcal{F})=\{\rho_{f}\;|\;f\in\mathcal{F}\}$. Then, the following evaluation holds.
@@
$\log\mathscr{N}(\delta,\psi\circ\rho(\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime})),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
@@
$\mathscr{N}(\delta,\psi\circ\rho(\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime})),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
$\leq\mathscr{N}(2^{-1}D_{\Delta^{d}}^{2}\delta,\rho(\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime})),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})}).$ (50)
@@
$\mathscr{N}(2^{-1}D_{\Delta^{d}}^{2}\delta,\rho(\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime})),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
@@
$\mathscr{N}(\delta,\psi\circ\rho(\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime})),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
@@
$\log\mathscr{N}(\delta,\psi\circ\rho(\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime})),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
@@
Note that the subclass $\mathcal{F}^{*}=\mathcal{F}_{L^{*},J^{*},S^{*},M^{*},\bm{d}^{*}}^{\Delta^{d}\textup{-NN}}$ is considered in this proof. Denote by $\mathcal{F}_{\textup{local}}^{*}=\mathscr{F}_{\beta,\beta^{-1}\varepsilon_{n},P}(\mathcal{F}^{*})$. Then, the set $\psi\circ\rho(\mathcal{F}_{\textup{local}}^{*})$ is $1$-uniformly bounded. Also, it is well known that hinge loss is $1$-Lipschitz continuous in the first argument (see, e.g., (Steinwart and Christmann, , 2008, Example 2.27)). Thus, we can set $c_{0}=1$ and $M_{n}=1$ hereafter.
@@
$\log\mathscr{B}(s,\psi\circ\rho(\mathcal{F}_{\textup{local}}^{*}),\|\cdot\|_{L^{2}(\mathcal{X}^{2})})\leq\log\mathscr{N}(\frac{s}{2},\psi\circ\rho(\mathcal{F}_{\textup{local}}^{*}),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})}).$ (63)
@@
$\log\mathscr{N}(2^{-1}s,\psi\circ\rho(\mathcal{F}_{\textup{local}}^{*}),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$